HUBBLE SPACE TELESCOPE DAILY REPORT #2283 PERIOD COVERED: 0000Z (UTC) 01/08/99 - 0000Z (UTC) 01/11/99 Daily Status Report as of 011/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Forty-Two Sets of NIC/1/2/3 8093 (DARKs: Warming Up - Part II) The Near Infrared Camera (Cameras One, Two and Three) was used to make a series of dark calibration observations. The purpose of this proposal is to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. (See 2.3) 1.3 Completed Two Sets of STIS/CCD/MA1/MA2 7482 (Long Slit Observations of the WR Central Star of the LMC-PN N66) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make long slit spectrographic observations of the Wolf-Rayet, central star of the planetary nebula LMC-N-66. We require long-slit STIS observations to cover from 1150 to 11000 Angstrom with a resolution of 1000 to 5000 so we can study the star and the nebula separately. A wavelength and a CCD flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Six Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.5 Completed Eleven Sets of NIC/1/2/3 8083 (FLATs: Warming Up - continuation) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. (See 2.3) 1.6 Completed WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxy 3C191.0. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena; they may be used as probes of their environments; and they are unique probes of the early Universe. The observations were completed as planned, and no problems were reported. 1.7 Completed FGS 7682 (FGS-3 Blue Transfer Function Reference Search Calibration) The Fine Guidance Sensor-3 was used to make observations of candidate star HD38666 in order to perform the blue Transfer Function reference search calibration. This proposal will measure the transfer function of a number of blue candidate stars in order to verify a single star which can be used for the blue calibration of the Trans Mode. The observations were exected as scheduled, and no anomalies were reported. 1.8 Completed Ten Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Five Sets of WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.10 Completed Ten Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation [[10[3I) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.11 Completed WF/PC-2 7377 (Dwarf Elliptical Galaxy Snapshot Survey II) The WF/PC-2 was used to make observations of the dwarf elliptical galaxy FCC128 as part of Snapshot Survey II. This survey will provide important information on the globular cluster systems of such galaxies and the properties of their cores and nuclei. The observations were completed as planned, and no anomalies were reported. 1.12 Completed Two Sets of NIC/2/3 8082 (NICMOS Pure Parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. (See 2.3) 1.13 Completed NIC/2/3 7962 (NICMOS Warm-up Focus Monitoring) The Near Infrared Camera (Cameras Two and Three) was used to make observations of the globular cluster and nebula of NGC-3603 as part of the NICMOS warm-up focus monitoring. This proposal is used to determine and monitor the optimal focus settings for NIC/2 and NIC/3 during the dewar warm-up period at the end of 1998. (See 2.3) 1.14 Completed WF/PC-2 6675 (The Evolution of Galaxies in Groups) The WF/PC-2 was used to make observations of the radio galaxy 5C12.251. There is now convincing evidence that galaxy evolution has progressed as a strong function of the environment. The strongest evolution is seen in the cores of rich clusters. The observations were completed as planned, and no anomalies were reported. 1.15 Completed Two Sets of WF/PC-2 6357 (The Nuclear Morphology of Elliptical Galaxies) The WF/PC-2 was used to make images of the early type, large, elliptical galaxies UGC-4551 and NGC-4589 (taken from a list of 150 galaxies) in order to study thier nuclear morphology. This survey will complement Illingworth's snapshot study of later Hubble types. The observations were completed as scheduled, and no anomalies were reported. 1.16 Completed STIS/CCD/MA1/MA2 7657 (STIS Spectroscopic and Imaging Sensitivity, MAMA) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of GD71 in order to study the spectroscopic and imaging sensitivity. Sensitivity measurements are done for all primary tilts of the gratings and the prism used with the MAMAs, except for the G140L and G230L gratings. Imaging sensitivity measurements are done with all apertures, including a neutral density filter. A wavelength and a dark calibration image were also taken. The loss of lock described in 2.1 and HSTAR 6884 occurred during this proposal. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.17 Completed Eight Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II parallel archive proposal continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.18 Completed STIS/CCD/MA1/MA2 7389 (Chromospheres and Transition Layers in Hyades and Pleiades F stars) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the F class star HD26345 which is located in the Hyades star cluster. This proposal seeks to obtain additional empirical constraints on the heating mechanisms for stellar chromospheres, transition layers and coronae. A wavelenth calibration obseration was also taken. The observations were completed as planned, and no problems were reported. 1.19 Completed WF/PC-2 6695 (The Stellar Populations in the LGS3 Dwarf Galaxy) The WF/PC-2 was used to make observations of the dwarf galaxy LGS3 in order to study its stellar populations. This galaxy is unique in that it has properties of both dwarf spheroidal and dwarf irregular systems, i.e. it is dominated by old or intermediate age stellar populations but it also contains a small amount of neutral hydrogen and a small young stellar population. The observations were completed as planned, and no anomalies were reported. 1.20 Completed Five Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.21 Completed NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. (See 2.3) 1.22 Completed WF/PC-2 6354 (Measuring Luminosity Evolution in Z = 0.3 Field Galaxies) The WF/PC-2 was used to make snapshot observations of the galactic cluster Gal-Cluster-1034 located at a distance of Z = 0.3. The targets were drawn from a much larger pool of galaxies for which the internal kinematics have been measured from an extensive ground based program. The observations were executed as scheduled, and no anomalies were noted. 1.23 Completed STIS/CCD/MA1/MA2 7753 (White Dwarf Companions of Barium and CH Peculiar Stars) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the white dwarf star HD65699-B which is a companion to a Barium and CH class star. In 1980 McClure et al.{ApJ 238, L35} discovered that all Ba stars show radial velocity variations indicative of their binary nature. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were noted. 1.24 Completed WF/PC-2 6671 (Deep V, I Photometry for 20 M31 Globular Clusters) The WF/PC-2 was used to make deep photometry observations (V=27.5) of the globular cluster M-31-G322 which is located in the Andromeda Galaxy (a.k.a., M-31). The observations were executed as planned, and no problems were reported. 1.25 Completed FGS 7597 (Long Term Stability of Transfer Scan Mode in FGS-1R) The Fine Guidance Sensors (FGS-1R) was used to make observations of the standard star UPGREN69 in order to study the long term stability of the Transfer Scan Mode. The observations were executed as scheduled, and no anomalies were noted. 1.26 Completed WF/PC-2 7387 (T-Tauri Star Snapshot Survey) The WF/PC-2 was used to make observations of the star DP-Tau as part of a T-Tauri class star survey. The observations were executed as scheduled, and no anomalies were noted. 1.27 Completed STIS/CCD 7438 (Spatially-Resolved Spectroscopy of the UVX in Elliptical Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to make spatially resolved far-ultraviolet observations of the large elliptical galaxy NGC-1399. The prominent far-UV upturn {UVX} in the spectra of elliptical galaxies originates from stars in their old populations which have evolved beyond the He-flash at the tip of the red giant branch. The observations were completed as scheduled, and no anomalies were reported. 1.28 Completed STIS/CCD/MA1 7473 (HST Observations of the DAB White Dwarf HS 0209+0832: Testing Explanations for the ''DB Gap'') The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make spectrographic observations of the DAB class white dwarf star HS0209-0832. White dwarfs are basically divided into two spectral classes: the hydrogen-rich DAs and the helium-rich non-DAs. However, between 28,000 and 45,000 K not a single non-DA star exists, although there are many helium-rich stars above and below this ``DB gap''. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 18 Successful : 18 There was a loss of lock at 009/015430Z. HSTAR 6884 was written. The proposal described in 1.16 may have been affected when the take data flag went down tempororily. Scheduled Reacquisitions : 33 Successful : 33 2.2 FHST Updates: Scheduled : 39 Successful : 36 As documented in HSTAR 6885, the full maneuver updates at 011/021244Z and at 011/021529Z failed. Subsequent acquisitions were successful, however. Per HSTAR 6886, the roll delay update at 011/033425Z, using FHSTs 1 and 2, failed. 2.3 Operations Notes: As noted in the 1/7/99 report, the NICMOS filter wheels have been inhibited. The proposals in this week's SMS (that ends with the period of this report) will execute. The instrument will transition to the SAA operate mode at the beginning of the day 011 SMS, and will remain in that mode until SM3. Using ROP SR-1, the SSR EDAC error counter was cleared nine times. A CCS/GRGT test was successfully concluded at 008/1851Z. The NSSC-1 status buffer was dumped at 008/1320Z and at 009/1328Z per ROP NS-3. The NICMOS FPA temperature sensors were enabled at 011/0010Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH